Hence, determine the probability of of a result with one die showing a 4 and the other a 5
1
Expert's answer
2020-10-11T17:35:27-0400

For the first die we need to get "4" while we can get any of {1,2,3,4,5,6} (6 elements in total). So the probability of this event is 1/6.

For the second die we need to get "5", and again there are 6 possible outcomes in total. The probability of showing "5" is again 1/6.

Since the events are independent ( the result of the second die doesn't depend of the result of the first die), we can multiply the probabilities:

"\displaystyle P(A) = \frac{1}{6} \cdot \frac{1}{6} = \frac{1}{36}"

Answer: 1/36.